OpenStudy (anonymous):

Mr Polaski is 5 times as old as his son David. Five years ago the father was 15 times as old as his son. What are their ages now?

OpenStudy (a_clan):

x=5y x-5=15(y-5)

OpenStudy (anonymous):

is the equation x-5=15(x-5)

OpenStudy (anonymous):

or is it 5x-5=15(x-5)

OpenStudy (anonymous):

let David's present age = x years then Polaski's age = 5x years five years ago Davidf's age = x-5 Polski's age = 5x -5 ATQ 5x -5 = 15(x-5) 5x - 5 = 15x -75

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

5x - 15x = -75 +5 -10x = -70 x = -70/-10 = 7

OpenStudy (anonymous):

so David is now 7 years Polaski is now 35 years

OpenStudy (anonymous):

ty so much wasn't sure why the equation was 5x-5=15(x-5) but got it now

OpenStudy (anonymous):

Good for you !!!! GOOD ANSWER click pls o(:})